OpenStudy (anonymous):

Rewrite f(x) = 3(x − 2)2 + 1 from vertex form to standard form. f(x) = 3x2 + 13 f(x) = 3x2 − 12x + 13 f(x) = 3x2 + 12x − 11 f(x) = 9x2 − 36x + 37

OpenStudy (anonymous):

@Nnesha

OpenStudy (arindameducationusc):

(x-2)^2 use (a-b)^2=a^2+b^2-2ab

OpenStudy (arindameducationusc):

in this case (x^2 +4-4x)

OpenStudy (arindameducationusc):

so it becomes 3(x^2 +4-4x)+1

OpenStudy (arindameducationusc):

just multiply now

OpenStudy (arindameducationusc):

it becomes.... Are you following me @boots_2000 ????

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is the answer d?

OpenStudy (arindameducationusc):

Nopes its not D

OpenStudy (arindameducationusc):

3(x^2 +4-4x)+1 =3x^2+12-12x+1 =3x^2-12x+13

OpenStudy (anonymous):

okay so its b right?

OpenStudy (arindameducationusc):

yes, but did you understand the steps?

OpenStudy (anonymous):

yes